# Disentangling representations of retinal images with generative models

### v1

@@ 4.2 Ablation study for disentanglement loss
We next conducted a sensitivity analysis for our prediction model with disentanglement loss and examined the effects of different hyperparameters. The subspace labels utilized for this analysis were age and camera. We specifically investigated the dimensionality of the subspace, the batch size, and the weight $\lambda_{DC}$ assigned to the loss term targeting the minimization of distance correlation.
@@
The initial hyperparameters were set as follows: a 4-dimensional age subspace $w_{\text{age}}$, a 12-dimensional camera subspace $w_{\text{cam}}$, a batch size of 512, and a loss weight of $\lambda_{DC}=0.5$ for distance correlation. Subsequently, each hyperparameter was individually varied, and for each configuration, we trained four models with different random initializations. All evaluation metrics were then reported on the test set. This analysis provided insight into the robustness and performance of our disentanglement loss under varying hyperparameters.
@@
Figure 8: Impact of hyperparameters on disentanglement performance. In a1-a3, we present performance variations across different subspace sizes. Specifically, a1 illustrates the distance correlation between subspaces, while a2 and a3 depict kNN accuracy improvement over chance level for age and camera, respectively. In b1-b3, we examine disentanglement performance under varying batch sizes. c1-c3 focus on the impact of different distance correlation loss weights on disentanglement performance. All reported values are derived from test data and four models with different initializations.
@@
The robustness of our disentanglement loss was evident across varied subspace sizes (Fig. 8a1-a3). Interestingly, there were no discernible trends when we altered the subspace size by multiplying our initial sizes with different factors. Notably, one exception was observed in the accuracy improvement for the camera within the age subspace, particularly when employing a multiplier of 4 (Fig. 8a3, blue curve). In this instance, the disentanglement problem becomes more challenging as we need to estimate the distance correlation between a 16- and a 48-dimensional subspace with a limited batch size of 512.
@@
Smaller batch sizes impacted the disentanglement performance (Fig. 8b1-b3). This effect was notable with the higher distance correlation values and worse decoding quality for a batch size of 8 compared to larger batch sizes (Fig. 8b1-b3). Notably, for the age subspace, a batch size of 64 proved insufficient, as evident by the camera decoding not reaching a possible minimum observed with larger batch sizes (blue curve, Fig. 8b3).
In the last ablation experiment, we explored the loss term $\lambda_{DC}$ for distance correlation minimization(Fig. 8c1-c3). The objective was to identify a valid range for this hyperparameter and to comprehend the consequences of deviating from this range. One non-surprising outcome was that the distance correlation became smaller with larger $\lambda_{DC}$ (Fig. 8c1). As expected, larger loss terms resulted in a collapse of subspace decoding. For example, the age accuracy improvement for both subspaces dropped below 10% (Fig. 8c2). Also, the disentanglement performance for the camera deteriorated with larger $\lambda_{DC}$, as the accuracy improvement in both subspaces converged to 30-40% (Fig. 8c3).
4.3 Learning disentangled subspaces with generative model
@@
Figure 9: Disentanglement performance with generative model for age-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{age}}$ encoding the age class $y_{\text{age}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, kNN classifier accuracy improvement over chance level accuracy is reported. The baseline model in a involved training linear classifiers solely on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces and our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@
As with the predictive model, as a baseline method, we learned subspaces by appending linear classifiers, addressing distinct classification tasks in each subspace. In the generative framework, we introduced an additional latent subspace dedicated to encoding the patient’s identity (Fig. 5), enhancing our ability to generate high-resolution and realistic retinal images. To ensure independence between these subspaces, we incorporated the minimization of distance correlation. Consequently, we aimed to minimize the distance correlation among three distinct subspace combinations. Notably, we specifically enforced independence for the patient-identity space, ensuring that it remained uncorrelated with the subspaces encoding age, ethnicity, or camera information.
@@
Figure 10: Disentanglement performance with generative model for ethnicity-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{ethn}}$ encoding the ethnicity class $y_{\text{ethn}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, kNN classifier accuracy improvement over chance level accuracy is reported. The baseline model in a involved training linear classifiers solely on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c illustrates the distance correlation measure between subspaces, comparing the baseline method to the model additionally minimizing distance correlation. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@ 4.4 Realistic image generation and reconstruction
Next, we evaluated the image quality and the reconstruction performance of our generative model. As a baseline we chose a standard StyleGAN2 model with GAN-inversion losses, which we compared to models incorporating our additional losses for independent subspace learning. The baseline model was trained with a 32-dimensional latent space and a single mapping network. For the comparison models, we chose our $\mathcal{W}$-spaces to be either 32- or 36-dimensional for the age-camera-identity and ethnicity-camera-identity setups, respectively. All models generated retinal fundus images with a resolution of $256\times 256$.
@@
Furthermore, we examined how additional subspace losses affected the image reconstruction performance, by reporting our inversion losses $L_{w}$ and $L_{p}$ for image encoding and pixel space reconstruction, respectively. As expected, image encoding $L_{w}$ consistently showed a slight degradation for models that we jointly optimized with our disentanglement loss (Tab. 1, model C and E). In contrast, the pixel space reconstruction loss $L_{p}$ remained comparable across all configurations. However, $L_{p}$ showed worse mean performance with high variances for some model configurations (Tab. 1, model A and D), suggesting that this loss could potentially benefit from a higher weight, which was also visible in the qualitative pixel space reconstruction difference maps (Fig. 11d-f).
@@
(D) Model C + subspace classifiers (ethnicity, camera) 12 $\pm$ 2 0.025 $\pm$ 0.01 0.010 $\pm$ 0.020
@@ 4.5 Controllable fundus image generation
In our final experiment, we explored how disentangled subspaces of a generative model influence the generation of retinal fundus images. We started by embedding the original images into latent spaces using our generative model. Because our models were trained with a disentanglement loss, were able to partition partitioned these latent spaces into three independent subspaces: the first subspace represented a patient attribute (age or ethnicity), the second subspace encoded the camera, and the third subspace represented patient-identity information. Subsequently, we performed subspace swapping by exchanging the subspace embeddings of different retinal images, allowing us to observe the effect of each subspace on image generation.
@@
We also provide quantitative evidence that the age subspace indeed influences image generation with relevant age features. To quantitatively assess the controllable image generation, we trained an image classifier on the task of predicting age from image reconstructions of our four GAN models. For detailed information about the training and testing procedures, please refer to Appendix 6.5. We tested the classifier models on a 3-class problem ($<50$, $50-60$, $\geq 60$) and a simpler 2-class problem distinguishing young from old patients ($\leq 40$, $\geq 65$). These two classification problems are almost balanced, with chance level accuracies of 37% for the 3-class problem and 57% for the 2-class problem. For testing, we evaluated the classification model in three scenarios: standard evaluation on test data, evaluation on reconstructed images from swapped age subspaces and their corresponding new age labels, and as a reference, we reconstructed images from swapped age subspaces but evaluated the classification performance against the original age labels. The third evaluation method ensured that the age information was not contained in the other subspaces. Comparing the three scenarios for both classification problems, we demonstrated that the age subspace indeed influenced image generation with the relevant features, as age predictions changed correctly when we swapped age subspaces. For example, in the 3-class problem, the performance dropped only from 70% to 66% for swapped age subspaces, compared to 40% for the wrong age label assignments (Tab. 2). Similarly, in the 2-class problem, we observed a drop from only 87% to 85%, compared to an accuracy of 47% for the wrong labels (Tab. 2).
@@ 5 Discussion
In this work, we focused on bridging the gap between existing generative models and disentangled representation learning for high-quality fundus image generation. We addressed the challenge of solving three joint tasks: (1) disentangling representations, (2) preventing shortcut learning, and (3) enabling controllable high-resolution image generation. To achieve this, we drew on recent advances in generative modeling and introduced a subspace GAN as a controllable population model for retinal fundus images. Our approach included the introduction of a novel disentanglement loss based on distance correlation. Through both qualitative and quantitative analyses, we successfully demonstrated the effectiveness of our disentangled subspaces within a generative model. Our approach is not only applicable to the field of medical imaging. The general prerequisites for our method can also be fulfilled in other areas: prior knowledge of confounding factors in the image generation process and labels for the primary task and the confounding factor.
@@
While our model shows excellent image generation performance with disentangled latent representations, some limitations should be acknowledged. One limitation is that the image reconstructions of our generative model showed deficits in capturing fine details, which is particularly evident in the accuracy of the vascular reconstructions. Interestingly, inherent correlations within the dataset pose a challenge for any disentanglement setup. Our experiments revealed that disentangling age from the camera was more straightforward than disentangling ethnicity from the camera, and that performance was limited by the strong correlation between ethnicity and camera in the EyePACS dataset (Träuble et al., 2021; Funke et al., 2022; Roth et al., 2023). In addition, it was challenging to disentangle the free identity space from the camera information. This difficulty may stem from the fact that these represent the largest subspace combinations, spanning 12- and 16-dimensional spaces, thus posing the greatest challenge for distance correlation estimation. Additionally, it is conceivable that camera features are correlated with other dataset features relevant to retinal fundus image generation, further complicating disentanglement.
@@
Moreover, our distance correlation-based disentanglement loss introduced some challenges. First, our model architecture made it necessary to find a suitable balance between the disentanglement loss and the image generation task, and over-weighting the disentanglement loss could lead to poor image generation performance. Second, our disentanglement loss also introduced some technical challenges. Notably, distance correlation computation is sensitive to batch size, necessitating the introduction of a ring buffer for multi-GPU training (Appendix 6.4). However, the size of the ring buffer was constrained by the need to balance stored batches from the past and a model that is updated every step. Furthermore, driven by distance correlation estimation, we used small latent space sizes that still enable high-quality retinal image generation. Consequently, distance correlation poses a limitation on scaling up the latent space. Larger latent space sizes would require correspondingly larger batch sizes for accurate distance correlation estimation. This interdependence prompts a deeper exploration of the dynamics between the data ring buffer and latent space size to enhance our understanding of their collective impact on model performance.
@@
In this study, our focus was exclusively on healthy fundus images due to the additional complexity that would have been introduced by reconstruction of disease features. However, the issue of spurious correlation between disease and technical features is a practical challenge that we aim to tackle in future work. Moreover, regardless of the fact that we deliberately chose to minimize the distance correlation as the disentanglement loss, alternative measures such as MMD and MI bounds or the use of adversarial classifiers can offer advantages, especially in terms of batch size or convergence time. We intend to investigate and compare these methods in the context of subspace disentanglement for medical images in future studies. In this work, we relied on labeled data to encode information into subspaces. Therefore, another compelling avenue for future research is to explore weakly supervised learning approaches, where we could apply a classification loss only on a subset of the training data, in addition to our disentanglement loss. This exploration could potentially yield effective results even on minimally labeled data.

### v2

@@ 4.2 Ablation study for disentanglement loss
Next, we performed a sensitivity analysis for our encoder model with disentanglement loss and investigated the effects of different hyperparameters. The subspace labels used for this analysis were age and camera. In particular, we examined the subspace dimensionality, the batch size, and $\lambda_{DC}$ (Eq. 3).
@@
The initial hyperparameters were set as follows: a 4-dimensional age subspace $w_{\text{age}}$, a 12-dimensional camera subspace $w_{\text{cam}}$, a batch size of 512, and $\lambda_{DC}=0.5$. Subsequently, each hyperparameter was individually varied, and for each configuration, we trained four models with different random initializations. All evaluation metrics were then reported on the test set.
@@
Figure 8: Impact of hyperparameters on disentanglement performance. In a1-a3, we present performance variations across different subspace sizes. Specifically, a1 illustrates the distance correlation between subspaces, while a2 and a3 depict the accuracy improvement over chance level for age and camera, respectively. In b1-b3, we examine disentanglement performance under varying batch sizes. c1-c3 focus on the impact of different distance correlation loss weights on disentanglement performance. We evaluated all metrics on the test data over four models with different initializations.
@@
The robustness of our disentanglement loss was evident across varied subspace sizes (Fig. 8a1-a3). Interestingly, there were no clear trends when we altered the subspace size by multiplying our initial sizes with different factors. Notably, one exception was observed in the accuracy improvement for the camera within the age subspace, particularly when employing a multiplier of 4 (Fig. 8a3, blue curve). In this case, the disentanglement problem becomes more challenging as we need to estimate the distance correlation between a 16- and a 48-dimensional subspace with a limited batch size of 512.
@@
Smaller batch sizes impacted the disentanglement performance (Fig. 8b1-b3). This effect was notable with the higher distance correlation values and worse decoding quality for a batch size of 8 compared to larger batch sizes (Fig. 8b1-b3). In particular, for the age subspace, a batch size of 8 and 64 proved to be insufficient, as shown by the camera decoding not reaching a possible minimum observed with larger batch sizes (Fig. 8b3, blue curve).
In our last ablation experiment, we explored the loss term $\lambda_{DC}$ for distance correlation minimization (Fig. 8c1-c3). As expected, the distance correlation measure between subspaces dropped with larger $\lambda_{DC}$ (Fig. 8c1). At the same time, the decoding accuracy for age from the age subspace (Fig. 8c2, blue curve) collapsed for $\lambda_{DC}\geq 2$, leading to a drop in decoding performance below 10%. In contrast, camera decoding (Fig. 8c2, green curve) from the age subspace becomes less accurate as soon as the disentanglement loss is introduced, indicating that the proposed method does what it is supposed to. For the camera subspace in (Fig. 8c3), the camera decoding accuracy remains high (green), and the age decoding accuracy initially drops (blue), in line with the goals of the distance correlation loss. However, the ablation study showed that the hyperparameter optimization of $\lambda_{DC}$ is essential, as over-weighting can lead to a complete collapse of the subspace encoding (Fig. 8c2,c3).
4.3 Learning disentangled subspaces with a generative model
@@
Figure 9: Disentanglement performance of a generative model for age-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{age}}$ encoding the age class $y_{\text{age}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@
As with the predictive model, as a baseline method, we learned subspaces by appending linear classifiers, addressing distinct classification tasks in each subspace. In the generative framework, we introduced an additional latent subspace dedicated to encoding the patient’s identity $w_{\text{id}}\in\mathbb{R}^{16}$ (Fig. 4), enhancing our ability to generate high-resolution and realistic retinal images. To ensure independence between these subspaces, we incorporated the minimization of distance correlation. Consequently, we aimed to minimize the distance correlation among three distinct subspace combinations. Notably, we specifically enforced independence for the patient-identity space, ensuring that it remained uncorrelated with the subspaces encoding age, ethnicity, or camera information.
@@
Figure 10: Disentanglement performance of a generative model for ethnicity-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{ethn}}$ encoding the ethnicity class $y_{\text{ethn}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@ 4.4 Realistic image generation and reconstruction
Next, we evaluated the image quality and the reconstruction performance of our generative model. As a baseline we chose a standard StyleGAN2 model with GAN-inversion losses, which we compared to models incorporating our additional losses for independent subspace learning. The baseline model was trained with a 32-dimensional latent space and a single mapping network. For the comparison models, we chose our $\mathcal{W}$-spaces to be either 32- or 36-dimensional for the age-camera-identity and ethnicity-camera-identity setups, respectively. All models generated retinal fundus images with a resolution of $256\times 256$ pixels.
@@
Furthermore, we examined how additional subspace losses affected the image reconstruction performance, by reporting our inversion losses $L_{w}$ and $L_{p}$ for image encoding and pixel space reconstruction, respectively. As expected, image encoding $L_{w}$ consistently showed a slight degradation for models that we jointly optimized with our disentanglement loss (Tab. 1, model C and E). In contrast, the pixel space reconstruction loss $L_{p}$ remained comparable across all configurations. However, $L_{p}$ showed worse mean performance with high variances for some model configurations (Tab. 1, model A and D), suggesting that this loss could potentially benefit from a higher weight, which was also visible in the qualitative reconstruction difference maps in pixel space (Fig. 11d-f).
@@
(D) Model A + subspace classifiers (ethnicity, camera) 12 $\pm$ 2 0.025 $\pm$ 0.01 0.010 $\pm$ 0.020
@@ 4.5 Controllable fundus image generation
In our final experiment, we explored how disentangled subspaces of a generative model influence the generation of retinal fundus images. We started by embedding the original images into latent spaces using our generative model. Because we trained our models with a disentanglement loss, we were able to split the latent spaces of our models into three independent subspaces: the first subspace represented a patient attribute (age or ethnicity), the second subspace encoded the camera, and the third subspace represented patient-identity information. Subsequently, we performed subspace swapping by exchanging the subspace embeddings of different retinal images, allowing us to observe the effect of each subspace on image generation.
@@
We also provide quantitative evidence that the age subspace indeed influences image generation with relevant age features. To quantitatively assess the controllable image generation, we trained an image classifier on the task of predicting age from image reconstructions of our four GAN models. For detailed information about the training and testing procedures, please refer to Appendix D. We tested the classifier models on a 3-class problem ($<50$, $50-60$, $\geq 60$) and a simpler 2-class problem distinguishing young from old patients ($\leq 40$, $\geq 65$). These two classification problems are almost balanced, with chance level accuracies of 37% for the 3-class problem and 57% for the 2-class problem. For testing, we evaluated the classification model in three scenarios: standard evaluation on test data, evaluation on reconstructed images from swapped age subspaces and their corresponding new age labels, and as a reference, we reconstructed images from swapped age subspaces but evaluated the classification performance against the original age labels. The third evaluation method ensured that the age information was not contained in the other subspaces. Comparing the three scenarios for both classification problems, we demonstrated that the age subspace indeed influenced image generation with the relevant features, as age predictions changed correctly when we swapped age subspaces. For example, in the 3-class problem, the performance dropped only from 70% to 66% for swapped age subspaces, compared to 40% for the wrong age label assignments (Tab. 2). Similarly, in the 2-class problem, we observed a drop from only 87% to 85%, compared to an accuracy of 47% for the wrong labels (Tab. 2).
@@ 5 Discussion
In this work, we focused on bridging the gap between existing generative models and disentangled representation learning for high-quality fundus image generation. We addressed the challenge of solving three joint tasks: (1) disentangling representations, (2) preventing shortcut learning, and (3) enabling controllable high-resolution image generation. To achieve this, we drew on recent advances in generative modeling and introduced a subspace GAN as a controllable population model for retinal fundus images. Our approach included the use of a disentanglement loss based on distance correlation. Through both qualitative and quantitative analyses, we successfully demonstrated the effectiveness of our disentangled subspaces within a generative model. Our approach is not only applicable to the field of medical imaging. The general prerequisites for our method can also be fulfilled in other areas: prior knowledge of confounding factors in the image generation process and labels for the primary task and the confounding factor.
@@
While our model shows excellent image generation performance with disentangled latent representations, some limitations should be acknowledged. One limitation is that the image reconstructions of our generative model showed deficits in capturing fine details, which is particularly evident in the accuracy of the vessel reconstructions. Interestingly, inherent correlations within the dataset pose a challenge for any disentanglement setup. Our experiments revealed that disentangling age from the camera was more straightforward than disentangling ethnicity from the camera, and that performance was limited by the strong correlation between ethnicity and camera in the EyePACS dataset (Träuble et al., 2021; Funke et al., 2022; Roth et al., 2023). In addition, it was challenging to disentangle the free identity space from the camera information. This difficulty may stem from the fact that these represent the largest subspace combinations, spanning 12- and 16-dimensional spaces, thus posing the greatest challenge for distance correlation estimation. Additionally, it is conceivable that camera features are correlated with other dataset features relevant to retinal fundus image generation, further complicating disentanglement. In general, approaches such as ours depend on the dataset distribution at hand and are therefore limited by inherent correlations. For many retrospective data sets obtained from routine care, patient demographics are determined by the visitors to the hospital. With our approach, we offer an algorithmic solution to disentangle (correlated) factors in the dataset to a certain degree.
@@
Moreover, our distance correlation-based disentanglement loss introduced some challenges. First, our model architecture made it necessary to find a suitable balance between the disentanglement loss, the feature encoding, and the image generation task, and over-weighting the disentanglement loss led to poor encoding and image generation performance. Second, our disentanglement loss also introduced some technical challenges. Notably, distance correlation computation is sensitive to the batch size as a hyperparameter, necessitating the introduction of a ring buffer for multi-GPU training (Appendix C). However, the size of the ring buffer was constrained by the need to balance stored batches from the past and a model that is updated every step. Furthermore, driven by distance correlation estimation, we used small latent space sizes that still enable high-quality retinal image generation. Consequently, distance correlation poses a limitation on scaling up the latent space. Larger latent space sizes would require correspondingly larger batch sizes for accurate distance correlation estimation. This interdependence prompts a deeper exploration of the dynamics between the data ring buffer, the latent space size, and the number of latent subspaces to improve our understanding of their joint impact on model performance.
To some extent, our method is scalable. Our method can be trained on larger datasets because the model training time scales linearly with the size of the dataset. In addition, our method can handle images of different resolutions well, as we show in the Appendix with Fig. F.3. Furthermore, our method can also handle combinations of multiple attributes well, as we show in Fig. G.4. Of course, the combinatorial complexity will eventually prohibit more attributes.
@@
In this study, our focus was exclusively on healthy fundus images due to the additional complexity that would have been introduced by the reconstruction of disease features. However, the issue of spurious correlation between disease and technical features is a practical challenge that we aim to tackle in future work. An additional latent disease subspace, e.g. learned with the EyePACS diabetic retinopathy (DR) labels, could be an interesting clinical test scenario for our work. Here, we could investigate whether a disentangled DR subspace is more robust for disease prediction on a shifted test distribution, e.g. from another hospital. However, in this study we only worked with healthy fundus images due to the unbalanced nature of the dataset. In another study (Ilanchezian et al., 2023), we used diffusion models for counterfactual image generation (without learning any representations), and disease features could be generated here with massive oversampling of the diseased classes, so this may be a way forward here as well. Moreover, alternative measures to distance correlation exist, such as maximum mean discrepancy (MMD) (Sejdinovic et al., 2013), MI bounds, or the use of adversarial classifiers, which can offer advantages, especially in terms of batch size or convergence time. We intend to investigate and compare these methods in the context of subspace disentanglement for medical images in future studies. In this work, we relied on labeled data to encode information into subspaces. Therefore, another compelling avenue for future research is to explore weakly supervised learning approaches, where we could apply a classification loss only on a subset of the training data, in addition to our disentanglement loss. This exploration could potentially yield effective results even on minimally labeled data. Lastly, the main contribution of this work is to extend an inverted GAN in general as an independent subspace learner that is applicable beyond the domain of fundus images. However, since fundus images have very specific properties (circular shape, fine vascular tree structure, optic disc, macula, etc.), biologically plausible biases could improve fundus image generation, which we leave for future work.